OpenStudy (anonymous):

Help please, will give medal. Given the two expressions shown below: A: 3 sqrt + 2 sqrt B: 3 sqrt + 6 sqrt Which statement best describes the two expressions? Choices: Both are rational. Both are irrational. A is rational, but B is irrational. A is irrational, but B is rational.

OpenStudy (anonymous):

3 sqrt is irrational Can you figure out the answer from there?

OpenStudy (anonymous):

You can add whatever you want to an irrational number, but it'll still be irrational

OpenStudy (anonymous):

The expressions: \[\sqrt{3} + \sqrt{2}\] \[\sqrt{3}+\sqrt{6}\]

OpenStudy (anonymous):

Both are irrational?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

You're welcome :]

OpenStudy (aum):

That's a rational answer :)